A small electric heating application

uses wire of 2 mm diameter with 0.8 mm thick insulation (k

= 0.12 W/m°C). The heat transfer coefficient (ho) on the in￾sulated surface is 35 W/m2°C. Determine the critical thick￾ness of insulation in this case and the percentage change in

the heat transfer rate if the critical thickness is used, assum￾ing the temperature difference between the surface of the

wire and surrounding air remains unchanged.




1
Expert's answer
2021-08-31T17:13:18-0400

Gives

"k=0.12W\/m\u00b0c"

"h_0=35W\/m^2\u00b0c"

"r_1=2mm\\r_2=1+0.8=2.8mm"

"\dot{Q_1}=\frac{∆t}{\frac{1}{2\pi kl}ln(\frac{r_2}{r_1})+\frac{1}{2\pi r_2l h_0}}"

"\dot{Q_1}=\frac{2\pi l∆t}{\frac{1}{k}ln(\frac{r_2}{r_1})+\frac{1}{r_2h_0}}"

Put value

"\dot{Q_1}=\frac{2\pi l∆t}{\frac{1}{0.12}ln(\frac{2.8}{2})+\frac{1}{0.0028\times35}}"

"\dot{Q_1}=\frac{2\pi l∆t}{0.02803+10.2040}"

"\dot{Q_1}=\frac{2\pi l∆t}{10.2321}"

"\dot{Q_1}=2\pi l∆t \times0.09773"

Critical radius


"r_c=\frac{k}{h_0}=\frac{0.12}{35}=0.0034285m=3.4285mm"

"\dot{Q_2}=\frac{2\pi l∆t}{\frac{1}{k}ln(\frac{r_c}{r_1})+\frac{1}{r_ch_0}}"

Put value

"\dot{Q_2}=\frac{2\pi l∆t}{\frac{1}{0.12}ln(\frac{3.4285}{2})+\frac{1}{0.0034285\times35}}"

"\dot{Q_2}=\frac{2\pi l∆t}{4.491+8.333}"

"\dot{Q_2}=\frac{2\pi l∆t}{12.8245}"

"\dot{Q_2}={2\pi l∆t}\times0.07797"

Percentage increase and heat loss


"=\frac{\dot{Q_2}-\dot{Q_1}}{\dot{Q_1}}\times100=\frac{0.07797-0.09773}{0.09773}\times100=20.21"

Heat loss=20.21%